OpenStudy (anonymous):

the volume of a sphere is 3000 pie . what is the surface area of the sphere to the nearest square meter? a.2158mm b. 524mm c. 314mm d.105mm

OpenStudy (blurbendy):

V = 4/3 times pi times r^3 V = 3000 pi 3000pi = 4/3pi * r^3 start by solving for r

OpenStudy (e.cociuba):

So your formula is formula is V=4/3*(pi*R^3)

OpenStudy (e.cociuba):

If v=3000 pi then, the Pi on both sides cancels out, so then 3000=4/3*r^3

OpenStudy (e.cociuba):

Multiply by reciprocal of 4/3 and get: 2250=r^3 Solve for radius (r)13.103707=r Now plug it in to the surface area formula which is sa=4*pi*r^2

OpenStudy (e.cociuba):

sa=4*pi*171.707137 sa=pi*686.828548 sa= 2157.73552 Round that up to 2158 for your answer of 2158 m^2 So your answers A.)